OpenStudy (anonymous):

Help me Simplify 9log9(7)

OpenStudy (owlcoffee):

\[\log_{a}B^n=nlog_{a} B\]

OpenStudy (anonymous):

9(log(9))(7) 60.117278 if this is wrong fell free to correct me

geerky42 (geerky42):

\(\log_ba = \dfrac{\ln a}{\ln b}\), I guess.

geerky42 (geerky42):

@aloud That's not simplifying though.

OpenStudy (anonymous):

I i think it is 9^7

geerky42 (geerky42):

I don't think you should move 9 to exponent in log. Doesn't seem like "simplifying" to me.

geerky42 (geerky42):

Not sure, To me, your expression is already simplified. Unless you can do something about base.

OpenStudy (anonymous):

after reworking it i get 7

OpenStudy (owlcoffee):

Neither is transforming into a neiperian logarithm with a denominator which complicates the domain.

geerky42 (geerky42):

"complicates the domain" Do you even know what domain is?

OpenStudy (anonymous):

Guys please i need help with this not argueing

OpenStudy (owlcoffee):

You can hardly simplify a logarithm, you are very far from lecturing me.

geerky42 (geerky42):

Tell me, genius. How does simplifying logarithm change its domain?

geerky42 (geerky42):

@ganeshie8

OpenStudy (anonymous):

Please help me......

geerky42 (geerky42):



OpenStudy (owlcoffee):

In spite of geerky, who clearly hasn't mastered basic algebra yet. We will use the property I stated above: \[\log_{a} B^n=nlog_{a} B\] So, for the problem: \[9\log_{9} (7)\] We'll just take it to the exponent: \[\log_{9}(7)^9 \] That would be a more simplified version, because later on, you'll learn that log(a)B^n is a more simplified way of expressing a logarithm because of the very definition: \[\log_{a}b^n=c <=>a^c=b^n\]

geerky42 (geerky42):

We are simplifying a expression, not equation. So turn it into exponential form wouldn't work out well. Plus it's not \(\log_9(7)^9\), it's \(\log_9(7^9)\)

geerky42 (geerky42):

Multiplication looks simplified to me than exponent. So as far as I concern, given expression is pretty much already simplified. Exactly what kind of form are you asked to simplify into? @maryah.g.osborn

OpenStudy (owlcoffee):

Do as you desire...

geerky42 (geerky42):

/thread

OpenStudy (owlcoffee):

hahahaha... As I said, do as you deem fit, I have no desire to argue with unmature people.

geerky42 (geerky42):

Clearly because you don't want to make yourself look any more fool.

ganeshie8 (ganeshie8):

relax friends i think the question looks terrible because of not using latex, i think it should be \[\large 9^{\log_9 7}\] @maryah.g.osborn please confirm if the expression looks as above or attach a screenshot if possible